# Control Pneumatic Soft Bending Actuator with Feedforward Hysteresis Compensation by Pneumatic Physical Reservoir Computing

### v1

@@ III-B Fuzzy PRC Hysteresis Compensation Model
It is essential to clarify that while the performance of the PI pressure controllers is assumed to be ideal in both the closed-loop control system and the FPRC model, these assumptions are grounded in different justifications. In the closed-loop control system, the PI controller is considered ideal primarily because the pressure control task is relatively straightforward and typically performs satisfactorily compared to the more complex task of controlling the soft actuator’s bending motion. Therefore, pressure control is not the main focus of this research. Conversely, the assumption of ideal performance for the PI controller in the FPRC model is based on the fact that the proportional valve, shown in Fig. 3 and controlled by the PI controller, also has its dynamics. The valve’s dynamics and any pressure control errors are integrated into the overall dynamics of the nonlinear physical system in the FPRC model. Therefore, the pressure control errors can be uncared.
@@
where $\tilde{P_{o}}$ is the filtered state of the physical reservoir. Unlike the state $\mathbf{x}$ in the ESN model, which can be defined with arbitrary dimensionality, the pneumatic physical reservoir’s state is represented by a single scalar $\tilde{P_{o}}$. Due to the absence of explicitly tunable hyperparameters, the adjustability of the physical system’s computational performance is comparatively limited. Our previous work used the time sequences of the physical reservoir’s output for state estimation [12]. Similarly, in this work, we collect the physical reservoir’s outputs (readouts) into temporal sequences of tap size $n_{u}$, as $\mathbf{p}(k)=[\tilde{P_{o}}(k),\tilde{P_{o}}(k-1),\cdots,\tilde{P_{o}}(k-n_{u}+1)]^{\top}\in\mathbb{R}^{n_{u}}$. The FPRC model’s performance can be adjusted by changing $n_{u}$.
Inspired by the precise hysteresis modeling achieved by using the T-S fuzzy model to process time sequences [14], we integrate a T-S fuzzy logic into the FPRC model. The T-S fuzzy model begins with clustering the extended state vector $\mathbf{x}^{\dagger}(k)=[\bm{\theta}^{\top}(k),\mathbf{p}^{\top}(k)]^{\top}\in\mathbb{R}^{n_{y}+n_{u}},\ k\in K_{\text{train}}$ into $n_{c}$ fuzzy clusters by the Fuzzy C-Means (FCM) algorithm [16], where $n_{c}$ is the number of T-S fuzzy rules. In each fuzzy cluster $i,\ i\in\{1,2,\cdots,n_{c}\}$, the state $\mathbf{x}^{\dagger}(k)$ has a membership $u_{i}(k)\in(0,1),\ \sum_{1}^{n_{c}}u_{i}(k)=1$, and is processed by an output layer $\mathbf{w}^{i}_{\text{out}}\in\mathbb{R}^{n_{y}+n_{u}+1}$ to produce an output $P^{i}_{\text{ff}}(k)$ as follows:
$P^{i}_{\text{ff}}(k)=\mathbf{w}^{i}_{\text{out}}\left[1,{\mathbf{x}^{\dagger}}^{\top}(k)\right]^{\top}=\mathbf{w}^{i}_{\text{out}}\left[\bm{\theta}^{\top}(k),\mathbf{p}^{\top}(k)\right]^{\top}.$ (7)
@@
$\mathbf{W}_{\text{output}}=\left[{\mathbf{w}^{1}}_{\text{output}}^{\top},{\mathbf{w}^{2}}_{\text{output}}^{\top},\cdots,{\mathbf{w}^{n_{c}}}_{\text{output}}^{\top}\right]^{\top}.$ (9)
During execution (including the model validation and test) the membership $\beta_{i}\in(0,1)$ of the state $\mathbf{x}^{\dagger}$ to the $i$-th fuzzy cluster is computed using the Gaussian membership function:
@@
where $\sigma$ is a scaling factor and $\mathbf{c}_{i}\in\mathbb{R}^{n_{y}+n_{u}}$ is the center of the $i$-th cluster obtained at the initial clustering using FCM. The obtained memberships are normalized into $\bar{\bm{\beta}}\in\mathbb{R}^{n_{c}}$:
@@ III-C Qualitative Model Comparison
The proposed FPRC model offers several advantages over the ESN model in the context of feedforward hysteresis compensation: 1) The FPRC model leverages the inherent nonlinear properties of physical systems, particularly the pneumatic soft actuator, to model hysteresis. Thus, it is less computationally demanding in real-time control applications than the ESN, as the physical system performs the nonlinear computation. 2) The FPRC model relies on a physical system for computation, where well-understood physical laws, e.g. the energy conservation, govern the physical system’s dynamics. Consequently, the responses of the physical system to specific input signals can be intuitively more understandable from a real-world view than that of the ESN’s hidden states, which may contain more than tens of dimensions [9].
Nevertheless, the FPRC model also presents several challenges: 1) The model’s reliance on an actually existing physical system introduces additional requirements for devices and space, such as the dual-PAM actuator, proportional valves, and pressure sensors used in this work, which potentially increases the system’s overall bulkiness. 2) Unlike the ESN, where performance can be easily tuned by hyperparameters, the computational capacity of the physical system is difficult to adjust and analyze. Additional techniques, e.g., the temporal sequence processing and the T-S fuzzy model employed in this work, may be necessary to compensate for this lack of flexibility in the PRC’s performance. These techniques may require domain expertise and introduce additional computational demands. 3) The exact mathematical description of the nonlinear dynamics, which is explicit in the ESN, can often be challenging to ascertain from the physical systems.
@@ III-D Quantitative Model Comparison
To quantitatively assess the ESN and FPRC models, we used the data shown in Fig. 4 (a), (b) and Fig. 5 (a) for model training, and used the data from Fig. 4 (c), (d) and Fig. 5 (b) to test the trained model. A comparative linear model (named Fuzzy linear) was constructed by substituting $[1,{\mathbf{x}^{\dagger}}^{\top}(k)]^{\top}$ in (7) and (12) with $[1,\bm{\theta}^{\top}(k)]^{\top}$ to show the necessity of the physical reservoir’s readout $\tilde{P_{o}}$ within the FPRC model.
The parameters for the ESN and FPRC models are listed in Table I. Before training and testing, a 100-step washout was employed to initialize the ESN’s state $\mathbf{x}$. The number of fuzzy rules $n_{c}$ in the comparative Fuzzy linear model was set to match that in the FPRC model. All models (i.e., FPRC, ESN, and Fuzzy linear) were configured with the tap size $n_{y}$ of 5 and the regularization factor $\alpha$ of 0.001. The parameter settings for the FPRC model were determined by trial and error, while those for the ESN were selected to align its performance with the FPRC model in the training phase.
TABLE I: Configuration Parameters for the Models
@@
All models underwent a 5-fold cross-validation training process, where each model was independently trained five times, each time using differently selected 80% of the training dataset for training and the remaining 20% for validation. The root-mean-square errors (RMSE) of the training ($e_{\text{train}}$) and the validation ($e_{\text{val}}$) were recorded during the cross-validation. The model yielding the minimal normalized error $\bar{e}=\sqrt{e^{2}_{\text{train}}+e^{2}_{\text{val}}}$ was deemed the optimal training outcome and subsequently tested. The training and test procedures were repeated ten times to compute the execution time costs.
Demonstrative training and test results of the FPRC model are presented in Fig. 8. The average and standard deviation of RMSE values during the 5-fold cross-validation, the RMSE values in the test phase, and the average and standard deviation of execution time are shown in Table II, where bold font highlights the best-performing model for each comparison metric. As shown in Fig. 8, the FPRC model’s results align closely with the ground truth in both the training and the test phase. Numerically, despite the similarities in training and validation performance between the ESN and FPRC models, the FPRC model demonstrates a lower RMSE in the test phase. The FPRC model’s test performance is consistent with its training performance, whereas the ESN model exhibits higher test RMSE values compared to the training and validation results. The RMSE comparison between the FPRC model and the Fuzzy linear model, where the only difference between them is the inclusion of $\tilde{P_{o}}$, highlights the contribution of the physical reservoir to the FPRC model’s improved accuracy. However, due to the additional features included, the FPRC model exhibits longer execution times than the Fuzzy linear model in training and testing. Compared to the ESN model, the FPRC model required less execution time for its training. Since the ESN’s state update, as (3) shows, involves the operation of a large matrix $\mathbf{W}_{r}$, which is instead processed by the physical system in the FPRC model, the FPRC model achieves a 97.6% reduction in execution time during the test phase compared to the ESN model.
Figure 8: Model training and test: (a) Training outcomes; (b) Test outcomes.
@@
We compared the three models by reversing the original training and test data for test and training (only here). Table III shows the results. Since the original training data contains more high-frequency signals than the original test data (as Fig. 8 shows), all models exhibit worse test RMSE than that in Table II due to unseen frequency features met at the test phase. The FPRC has the lowest RMSEs among all models, highlighting its generalization with unseen information.
@@
The higher magnitude of the linear weight can indicate its corresponding nonlinear feature’s more predominant role in describing the nonlinear dynamics [17]. To quantitatively clarify the physical reservoir’s contribution within the FPRC model, we normalized the training data as follows:
@@
where $x$ represents either $\tilde{P_{o}}$ or $\theta$ (used as $\theta_{d}$ in training), and $\min(x)$ ($\max(x)$) denotes the minimum (maximum) value of $x$ within the training dataset. The normalized data were then used to train the FPRC model by the aforementioned procedure, and the obtained weights were normalized by:
@@
where ||$\cdot$|| ⋅ | is the absolute value, and $\omega_{j},\ j\in\{1,2,\cdots,n_{u}+n_{y}+1\}$ is the individual linear weight. The normalized weights corresponding to different nonlinear features of the FPRC model, including $\tilde{P_{o}}$, $\theta_{d}$, and the bias term, are shown in Fig. 9. As shown in Fig. 9, the importance of different features varies among different fuzzy rules. The upper dashed red line highlights that the weights combining $\tilde{P_{o}}$ exhibit the same magnitude as those processing $\theta_{d}$. On average, the weights corresponding to $\tilde{P_{o}}$ and $\theta_{d}$ account for 56.1% and 43.8%, respectively, across the 8 fuzzy rules.
@@ III-E Performance Analysis Based on Control Variable
We used the control variable method to analyze the effect of different parameter settings on the FPRC model’s performance. Specifically, we used the filter factor $\epsilon$ being 1, 0.1, 0.01, 10-3, and 10-4 to train the FPRC model with all other parameters consistent with the previous section. Similarly, we trained the FPRC and the Fuzzy linear models with $n_{c}$ being 1, 2, 4, 8, and 16, with all other parameters fixed. The results are presented in Fig. 10. The results show that the increase in $\epsilon$ and $n_{c}$ causes the FRPC model’s training RMSE to decrease, while the test RMSE exhibits complex non-monotonic variations. Results in (c) show that greater $n_{c}$ leads to longer training time as a larger group of weights are learned, but the test time cost barely changes. Despite the longer execution times, the FPRC model consistently exhibits lower training, validation, and test RMSE values than the Fuzzy linear model with all tested parameter settings.
Figure 10: Training and test results: (a) Variation of RMSEs with $\epsilon$ values; (b) and (c) Variation of RMSEs and execution time costs with $n_{c}$ values.
The previous section has illustrated the significance of including $\tilde{P_{o}}$ in the FPRC model (i.e., $n_{u}>0$). We extended this investigation by training the FPRC model with different $n_{u}$ and $n_{y}$ combinations (ranging from 1 to 10) with all other parameters fixed, and the results are shown in Fig. 11. It is observed that while the model’s training and validation performance benefits from less $n_{u}$ and greater $n_{y}$, a medium $n_{u}$ helps decrease the RMSE value in the model test stage.
Figure 11: Model errors with different tap sizes, including RMSEs in training, validation, and test stage. Results were obtained with $n_{c}=8$, and $\epsilon=0.01$.
Theoretically explaining the test error’s variation with model parameters, e.g., the ‘double descent’ phenomenon [18], is difficult and needs domain expertise. As our scope is developing the FPRC model to control soft actuators, further analysis of the test error’s complex variation is omitted here.
@@ IV Experimental Results and Discussion
Experimental results of tracking sine waves of 0.2Hz and 0.5Hz are shown in Fig. 12 and Fig. 13. The results indicate that both the open-loop control using only the FPRC model and the closed-loop FPRC+PD system closely followed the reference signals.
Figure 12: Tracking a 0.2Hz sine wave: (a) Open-loop control; (b) Closed-loop control; (c) Errors; (d) Feedforward and feedback control outputs.
The closed-loop control method performs better because of the feedback controller’s adjustment capabilities. Variations of the feedforward model’s output $P_{\text{ff}}$ and the feedback adjustments $P_{\text{fb}}$ during the experiments are shown in Fig. 12(d) and Fig. 13(d). These results demonstrate that the FPRC model’s output $P_{\text{ff}}$ is the predominant element in the closed-loop control system, with the PD controller’s minimal output $P_{\text{fb}}$ mainly facilitating fine adjustments.
Figure 13: Tracking a 0.5Hz sine wave: (a) Open-loop control; (b) Closed-loop control; (c) Errors; (d) Feedforward and feedback control outputs.
Additional experiments were conducted by tracking a chirp signal and a complex signal, with results shown in Fig. 14 and Fig. 15. Results from Fig. 14 show that although the tracking error increases with the reference signal’s frequency, it remains within ±$\pm 5$ [deg] in the closed-loop system without detectable phase lag at high-frequency zones. Results in Fig. 15 exhibit the FPRC model’s capability to control the actuator’s bending following a varying signal. The numerical comparison in tracking RMSEs [deg] is shown in Table IV.
Figure 14: Tracking a sweep-frequency chirp signal: (a) Open-loop control; (b) Closed-loop control; (c) Errors; (d) Feedforward and feedback control outputs. The desired bending angle during the experiment is $32.527.5\sin\left(\pi t(0.01125t+0.1)-0.5\pi\right)+$ 32.5 [deg], $t\in[0,80]$ is the time [s].
Figure 15: Tracking a complex signal: (a) Open-loop control; (b) Closed-loop control; (c) Errors; (d) Feedforward and feedback control outputs. The reference signal is $40.56.5\sum_{i=1}^{5}\sin\left(2\pi f_{i}t-\pi/2\right)+$ 40.5 [deg], $t\in[0,100]$ is the time [s], with $f_{1}=0.12$, $f_{2}=0.04$, $f_{3}=0.31$, $f_{4}=0.29$, and $f_{5}=0.25$.
@@
Control system 0.2 Hz sine 0.5 Hz sine Chirp Complex
FPRC 2.04 2.25 2.67 1.70
FPRC+PD 0.795 1.16 1.70 0.608
To demonstrate the FPRC+PD method’s superior performance over the standalone PD controller, we utilized the PD controller to track the signals shown in Fig. 13 and Fig. 14 and compared its performance with the FPRC+PD system, with the results shown in Fig. 16. The PD controller exhibits noticeable hysteresis loops in the reference-to-actual relationship. In contrast, including the FPRC feedforward model effectively mitigates such hysteresis loops.
Figure 16: Control results comparison between FPRC+PD and PD: (a) Track the reference signal in Fig. 13; (b) Track the reference signal in Fig. 14.
A robustness test was conducted using the FPRC+PD to control the actuator’s bending following a 0.3 Hz sine wave. Initially, the pneumatic physical reservoir operated without disturbances for a short period. Subsequently, the system was subjected to manually imposed random disturbances. A bending sensor recorded the physical reservoir’s bending state during the experiment to provide an intuitive visualization of the disturbance’s effects on the physical reservoir. The tracking results and the physical reservoir’s bending angle are presented in Fig. 17 (a) and (b), respectively. The experimental results in (b) exhibit that while the physical reservoir’s bending state was significantly affected by external disturbances, its output $P_{o}$ remained stable and was not substantially influenced. The tracking results in (a) further indicate that the control system’s performance was unaffected by the disturbances applied to the physical reservoir.
The physical reservoir’s exceptional environmental robustness can be attributed to the fact that variations in the $P_{o}$ are caused by the compression of the inextensible fabric shield, which is driven by the inflation of the active PAM rather than by the reservoir’s bending. Consequently, regardless of the reservoir’s bending being disturbed, the pressurization $P_{i}$ of the active PAM compels the fabric to compress the passive PAM, ensuring the physical reservoir’s robust operation.
Figure 17: Tracking a 0.3 Hz sine signal with external disturbance: (a) Closed-loop control tracking result; (b) Input output pressures and the bending state of the pneumatic physical reservoir. Results are depicted with background colors of yellow for no disturbance and gray for under disturbances.

### v2

@@ III-B Fuzzy PRC Feedforward Hysteresis Compensation Model
where $\tilde{P_{o}}$ is the filtered state of the physical reservoir. Unlike the state $\mathbf{x}$ in the ESN, which can be defined with arbitrary dimensionality, the physical reservoir’s state is represented by a single scalar $\tilde{P_{o}}$. Due to the absence of explicitly tunable hyperparameters, the adjustability of the physical reservoir’s computational performance is comparatively limited [16]. Previous works used the time sequences of the physical reservoir’s readouts for computation [12, 10]. Similarly, we collect the physical reservoir’s outputs into temporal sequences of size $n_{u}$, as $\mathbf{p}(k)=[\tilde{P_{o}}(k),\tilde{P_{o}}(k-1),\cdots,\tilde{P_{o}}(k-n_{u}+1)]^{\top}\in\mathbb{R}^{n_{u}}$. The FPRC model’s properties can then be adjusted by altering $n_{u}$.
Motivated by the precise hysteresis modeling achieved by using the T-S fuzzy model to process time sequences [14], we integrate a T-S fuzzy logic into the FPRC model. The T-S fuzzy model begins with clustering the extended state vector $\mathbf{x}^{\dagger}(k)=[\bm{\theta}^{\top}(k),\mathbf{p}^{\top}(k)]^{\top}\in\mathbb{R}^{n_{y}+n_{u}},\ k\in K_{\text{train}}$ into $n_{c}$ fuzzy clusters by the Fuzzy C-Means (FCM) algorithm [17], where $n_{c}$ is the number of T-S fuzzy rules. In each fuzzy cluster $i,\ i\in\{1,2,\cdots,n_{c}\}$, the state $\mathbf{x}^{\dagger}(k)$ has a membership $u_{i}(k)\in(0,1)\ s.t.\sum_{1}^{n_{c}}u_{i}(k)=1$, and is processed by an output layer $\mathbf{w}^{i}_{\text{out}}\in\mathbb{R}^{n_{y}+n_{u}+1}$ to produce an output $P^{i}_{\text{ff}}(k)$ as follows:
$P^{i}_{\text{ff}}(k)=\mathbf{w}^{i}_{\text{out}}\left[1,{\mathbf{x}^{\dagger}}^{\top}(k)\right]^{\top}=\mathbf{w}^{i}_{\text{out}}\left[1,\bm{\theta}^{\top}(k),\mathbf{p}^{\top}(k)\right]^{\top}.$ (7)
@@
$\mathbf{W}_{\text{out}}=\left[{\mathbf{w}^{1}}_{\text{out}}^{\top},{\mathbf{w}^{2}}_{\text{out}}^{\top},\cdots,{\mathbf{w}^{n_{c}}}_{\text{out}}^{\top}\right]^{\top}.$ (9)
During execution (including the model validation and test) the membership $\beta_{i}\in(0,1)$ of the state $\mathbf{x}^{\dagger}$ to the $i$-th fuzzy cluster is computed using the Gaussian membership function as follows:
@@
where $\sigma$ is a scaling factor and $\mathbf{c}_{i}\in\mathbb{R}^{n_{y}+n_{u}}$ is the center of the $i$-th cluster obtained at the initial clustering using FCM. The obtained memberships are normalized into $\bar{\bm{\beta}}\in\mathbb{R}^{n_{c}}$ as:
@@ III-C Qualitative Model Comparison
The FPRC model offers several advantages over the ESN model regarding the implementation as the feedforward hysteresis compensation model: 1) The FPRC model leverages the inherent nonlinear properties of physical systems, precisely, the dual-PAM soft actuator in this work, to model hysteresis. Hence, it is less computationally expensive in real-time control applications than the ESN, as the real-world nonlinear physical system performs the state update [10]. 2) The FPRC model relies on a real-world physical system for computation, where well-understood physical laws, e.g., energy conservation, govern the physical system’s dynamical behaviors. Consequently, the physical system’s responses to specific input signals are usually monotonic and thus intuitively more understandable from a real-world view than the ESN’s complex state variations that may contain more than tens of dimensions [9].
Nevertheless, the FPRC model also presents several disadvantages: 1) The model’s reliance on an actually existing physical system introduces additional requirements for devices and space, such as the dual-PAM actuator, proportional valves, and pressure sensors used in this work, which potentially increases the system’s overall bulkiness. 2) Unlike the ESN, where performance can be easily tuned by hyperparameters, the computational capacity of the physical system is difficult to adjust and analyze [16]. Additional techniques, e.g., the temporal sequence processing and the T-S fuzzy model employed in this work, may be necessary to compensate for this lack of flexibility in the PRC’s performance. These techniques may require domain expertise and introduce additional computational demands. 3) The exact mathematical description of the nonlinear dynamics, which is explicit in the ESN, can often be challenging to ascertain from real-world physical systems.
@@ III-D Quantitative Model Comparison
To quantitatively assess the ESN and FPRC models, we used the data shown in Fig. 4 (a), (b) and Fig. 5 (a) for model training, and used the data from Fig. 4 (c), (d) and Fig. 5 (b) to test the trained model. A comparative linear model (named Fuzzy linear) was constructed by substituting $[1,{\mathbf{x}^{\dagger}}^{\top}(k)]^{\top}$ in (7) and (12) with $[1,\bm{\theta}^{\top}(k)]^{\top}$ to illustrate the contribution of the physical reservoir’s readout $\tilde{P_{o}}$ within the FPRC model.
Table I lists the parameters for ESN and FPRC models. Before model training and test, a 100-step washout was used to initialize the ESN’s state $\mathbf{x}$. The number of fuzzy rules $n_{c}$ in the comparative Fuzzy linear model was set consistent with the FPRC model. All models (i.e., FPRC, ESN, and Fuzzy linear) have the tap size $n_{y}$ of 5 and the ridge parameter $\alpha$ of 0.001. The parameters for the FPRC model were chosen by trial and error, while those for the ESN were selected to align its performance with the FPRC model in the training stage.
TABLE I: Configuration Parameters for Different Models
@@
All models were trained by a 5-fold cross-validation process, where each model was independently trained five times, each time using differently selected 80% of the training set for training and the remaining 20% for validation. The root-mean-square errors (RMSE) of the training ($e_{\text{train}}$) and the validation ($e_{\text{val}}$) were recorded during the cross-validation. The model yielding the minimal normalized error $\bar{e}=\sqrt{e^{2}_{\text{train}}+e^{2}_{\text{val}}}$ was deemed the optimal training outcome and subsequently evaluated on the test dataset. The training and test procedures were repeated ten times to compute the execution time costs.
Demonstrative training and test results of the FPRC model are presented in Fig. 8. The average and standard deviation of RMSE values during the 5-fold cross-validation, the RMSE values in the test phase, and the average and standard deviation of execution time are shown in Table II, where bold font highlights the best-performing model for each comparison metric. As shown in Fig. 8, the FPRC model’s results align closely with the ground truth $P_{\text{exp}}$ in both the training and the test phase. Numerically, despite the similarities in training and validation performance between the ESN and FPRC models, the FPRC model demonstrates a lower RMSE in the test phase. The FPRC model’s test RMSE is consistent with its training outcomes, whereas the ESN model exhibits higher test RMSE values compared to its training and validation results. The RMSE comparison between the FPRC model and the Fuzzy linear model, where the only difference between them is the inclusion of $\tilde{P_{o}}$, highlights the contribution of the physical reservoir to the FPRC model’s improved accuracy. However, due to the additional features ($\tilde{P_{o}}$) included, the FPRC model exhibits longer execution times than the Fuzzy linear model. Compared to the ESN model, the FPRC model required less execution time for its training. Since the ESN’s state update, as (3) shows, involves the operation of a large matrix $\mathbf{W}_{r}$, which is instead processed by the physical system in the FPRC model, the FPRC model achieves a 97.6% reduction in execution time during the test phase than the ESN model.
Figure 8: Training and test results of the FPRC model: (a) Training; (b) Test.
@@
We compared the three models by reversing the original training and test data for test and training (only here). Table III shows the results. Since the original training data contains more high-frequency signals than the original test data (as Fig. 8 shows), all models exhibit worse test RMSE than that in Table II due to unseen frequency features met at the test phase. The FPRC exhibits the lowest RMSEs among all models, highlighting its better generalization with unseen information.
@@
The greater magnitude of a linear weight can indicate its corresponding nonlinear feature’s more predominant position in describing the nonlinear dynamics [18]. To quantitatively clarify the physical reservoir’s contribution within the FPRC model, we first normalized the training data by the following:
@@
where $x$ represents either $\tilde{P_{o}}$ or $\theta$ (used as $\theta_{d}$ in training), and $\min(x)$ ($\max(x)$) denotes the minimum (maximum) value of $x$ within the training dataset. Then, the normalized data were used to train the FPRC model by the aforementioned process, and the obtained linear weights were normalized as follows:
@@
where ||$\cdot$|| ⋅ | means the absolute value, and $\omega_{j},\ j\in\{1,2,\cdots,n_{u}+n_{y}+1\}$ denotes the individual linear weight (including bias) contained in $\mathbf{w}^{q}_{\text{out}}$ of each fuzzy rule $q$, $q\in\{1,2,\cdots,n_{c}\}$. The normalized weights corresponding to different nonlinear features of the FPRC model, including $\tilde{P_{o}}$, $\theta_{d}$, and the bias term, are shown in Fig. 9. As shown in Fig. 9, the importance of different features varies among different fuzzy rules. The upper dashed red line highlights that the weights combining $\tilde{P_{o}}$ exhibit the same magnitude as those processing $\theta_{d}$. On average, the weights corresponding to $\tilde{P_{o}}$ and $\theta_{d}$ account for 56.1% and 43.8% across the eight fuzzy rules, respectively.
@@ III-E Performance Analysis Based on Control Variable
We used the control variable method to analyze the influence of different parameters on the FPRC model’s performance. Specifically, we used the filter factor $\epsilon$ being 1, 0.1, 0.01, 10-3, and 10-4 to train the FPRC model with all other parameters consistent with the previous section. Similarly, we trained the FPRC and the Fuzzy linear models with $n_{c}$ being 1, 2, 4, 8, and 16, with all other parameters fixed. The results are presented in Fig. 10. The results show that the increase in $\epsilon$ and $n_{c}$ causes the FRPC model’s training RMSE to decrease, while the test RMSE exhibits complex non-monotonic variations. Results in (c) show that greater $n_{c}$ leads to longer training time as a larger group of weights are optimized, but the test time cost barely changes. Despite the longer execution times, the FPRC model consistently shows lower training, validation, and test RMSE values than the Fuzzy linear model across the tested settings.
Figure 10: Training and test results: (a) Variations of RMSEs with $\epsilon$ values; (b) and (c) Variations of RMSEs and execution time costs with $n_{c}$ values.
The previous section has shown the necessity of $\tilde{P_{o}}$ within the FPRC model, i.e., $n_{u}>0$. We extended this analysis by training the FPRC model with different $n_{u}$ and $n_{y}$ (from 1 to 10) combinations with all other parameters fixed, and the results are shown in Fig. 11. It is observed that while the model’s training and validation performance benefits from less $n_{u}$ and greater $n_{y}$, a medium $n_{u}$ selection helps decrease the RMSE value in the model test stage, benefiting its generalization.
Figure 11: Model errors with different tap sizes, including RMSEs in training, validation, and test stages. Results were obtained with $n_{c}=8$, and $\epsilon=0.01$.
Theoretically explaining the test performance’s variation with model parameters, e.g., the ‘double descent’ phenomenon [19], is challenging and needs domain expertise. As our scope is developing the FPRC model to control soft actuators, further analysis of the test error’s complex variation is omitted here.
@@ IV Experimental Results and Discussion
Experimental results of tracking 0.2 Hz and 0.5 Hz sine waves are shown in Fig. 12 and Fig. 13. The results indicate that both the open-loop control with the standalone FPRC model and the closed-loop FPRC+PD control system closely followed the reference signals. The closed-loop control method performs better because of the feedback controller’s adjustment capabilities. Variations of the feedforward model’s output $P_{\text{ff}}$ and the feedback adjustments $P_{\text{fb}}$ during experiments are shown in Fig. 12(d) and Fig. 13(d). These results show that the FPRC model’s output $P_{\text{ff}}$ is the predominant element in the FPRC+PD closed-loop control system, with the PD controller’s output $P_{\text{fb}}$ primarily serving as fine adjustments.
Figure 12: Tracking a 0.2Hz sine wave: (a) Open-loop control; (b) Closed-loop control; (c) Control errors; (d) Feedforward and feedback controller outputs.
Figure 13: Tracking a 0.5Hz sine wave: (a) Open-loop control; (b) Closed-loop control; (c) Control errors; (d) Feedforward and feedback controller outputs.
Additional experiments were conducted by tracking a chirp signal and a complex signal, with results shown in Fig. 14 and Fig. 15. Results from Fig. 14 show that although the bending-motion tracking error increases with the desired signal’s frequency, it remains within ±$\pm 8$ [deg] in the closed-loop system without detectable phase lag at high-frequency zones. Results in Fig. 15 exhibit the FPRC model’s capability to control the actuator’s bending following a varying signal. The numerical comparison of tracking errors in RMSE is shown in Table IV.
Figure 14: Tracking a sweep-frequency reference signal: (a) Open-loop control; (b) Closed-loop control; (c) Control errors; (d) Feedforward and feedback controller outputs. The reference bending angle is expressed as $32.527.5\sin\left(\pi t(0.01125t+0.1)-0.5\pi\right)+$ 32.5 [deg], $t\in[0,80]$ is the time [s].
Figure 15: Tracking a complex signal: (a) Open-loop control; (b) Closed-loop control; (c) Control errors; (d) Feedforward and feedback control outputs. The reference signal is $40.56.5\sum_{i=1}^{5}\sin\left(2\pi f_{i}t-\pi/2\right)+$ 40.5 [deg], where $t\in[0,100]$ is the time [s], with $f_{1}=0.12$, $f_{2}=0.04$, $f_{3}=0.31$, $f_{4}=0.29$, and $f_{5}=0.25$.
@@
Control system Fig. 12 Fig. 13 Fig. 14 Fig. 15
FPRC 2.04 [deg] 2.25 [deg] 2.67 [deg] 1.70 [deg]
FPRC+PD 0.795 [deg] 1.16 [deg] 1.70 [deg] 0.608 [deg]
To demonstrate the FPRC+PD control system’s superior performance over the standalone PD controller, we used the PD controller to track the signals shown in Fig. 13 and Fig. 14 and compared its tracking outcomes with that from the FPRC+PD system. The results are shown in Fig. 16, where the PD controller exhibits noticeable hysteresis loops in the reference-to-actual relationship. In contrast, including the FPRC feedforward model largely mitigates such loops.
Figure 16: Control results comparison between FPRC+PD and PD: (a) Tracking the reference signal in Fig. 13; (b) Tracking the reference signal in Fig. 14.
A robustness test was conducted using the FPRC+PD to control the actuator’s bending following a 0.3 Hz sine wave. Initially, the pneumatic physical reservoir worked without disturbances for a short period. Subsequently, the physical reservoir was subjected to manually imposed random disturbances. A bending sensor recorded the physical reservoir’s bending motion during the experiment to provide an intuitive vision of the disturbance’s effects on the physical reservoir. The tracking results and the physical reservoir’s bending angle are shown in Fig. 17 (a) and (b). The experimental results in (b) exhibit that while the physical reservoir’s bending was significantly distorted by disturbances, its output $P_{o}$ remained stable and was not substantially influenced. The tracking results in (a) further indicate that the control system’s performance was unaffected by the disturbances applied to the physical reservoir.
The physical reservoir’s exceptional environmental robustness can be attributed to the fact that variations in $P_{o}$ are caused by the compression of the inextensible fabric shield, which is primarily driven by the inflation of the active PAM rather than by the reservoir’s actual bending. Consequently, regardless of the reservoir’s bending being disturbed, the pressurization $P_{i}$ of the active PAM compels the fabric to squeeze the passive PAM, thus ensuring the physical reservoir’s robust operation.
Figure 17: Tracking a 0.3 Hz sine signal with external disturbance: (a) Closed-loop control tracking result of the actuator; (b) Input and output pressures and bending state of the physical reservoir. Results are depicted with background colors of yellow for no disturbance and gray for under manual disturbances.